OpenStudy (anonymous):

complete the square, and find the roots of the quadratic equation. 5. (4 points) x2 + 6x -16= 0

OpenStudy (katrinakaif):

x = -2 or x = 8

OpenStudy (anonymous):

complete the square like this: \[x^2+6x-16=0\] \[x^2+6x=16\] \[(x+3)^2=16+9\] \[(x+3)^2=25\] \[x+3=5\] or \[x+3=-5\] \[x=2\] or \[x=-8\]